# Advancing Pose-Guided Image Synthesis with Progressive Conditional Diffusion Models

### v1

@@ Appendix B Implement Details
Our experiments are conducted on 8 NVIDIA V100 GPUs. We utilize the AdamW optimizer with a consistent learning rate of $1e^{-4}$ across all stages. The probability of random dropout for condition $c$ is set at 10%. For the prior conditional diffusion model, we employ OpenCLIP ViT-H/14  [^2] https://github.com/mlfoundations/open_clip as the CLIP image encoder. This model’s transformer consists of 20 transformer blocks, each with a width of 2,048. The model is trained for 10k iterations with a batch size of 256, using a cosine noising schedule (Nichol & Dhariwal, 2021) with 1000 timesteps. For the inpainting and refining models, we use DINOv2-G/14  [^3] https://github.com/facebookresearch/dinov2 as the image encoder. We leverage the pretrained stable diffusion V2.1  [^4] https://huggingface.co/stabilityai/stable-diffusion-2-1-base, modifying the first convolution layer to accommodate additional conditions. These models are trained for 200k and 100k iterations, respectively, each with a batch size of 128, and a linear noise schedule with 1000 timesteps is applied. In the inference stage, we use the DDIM sampler with 50 steps and set the guidance scale w to 2.0 for PCDMs.
@@ Appendix C Additional Results
C.1 More Qualitative Comparisons for PCDMs
We provide additional examples for comparison with the state-of-the-art (SOTA) methods in Figure 10 and Figure 11.
@@ C.1 More Qualitative Comparisons for PCDMs
C.2 More qualitative ablation results on the DeepFashion dataset
@@
C.3 Random Samples
We show random samples from our proposed PCDMs in Figure 13. The generated results are highly stable and realistic.
Figure 13: Random samples from PCDMs.
C.4 User Study

### v3

@@ Appendix B Implement Details
Table 4: Hyperparameters for the PCDMs. All models trained on 8 V100 GPUs.
Hyperparameters Frist Stage Second Stage Third Stage
Diffusion Steps 1000 1000 1000
Noise Schedule cosine linear linear
Optimizer AdamW AdamW AdamW
Weight Decay 0.01 0.01 0.01
Batch Size 256 128 128
Iterations 100k 200k 100k
Learning Rate $41e-$ 4 $41e-$ 4 $41e-$ 4
Our experiments are conducted on 8 NVIDIA V100 GPUs. We follow the standard training strategies and hyperparameters of diffusion models. We utilize the AdamW optimizer with a consistent learning rate of $1e^{-4}$ across all stages. The probability of random dropout for condition $c$ is set at 10%. For the prior conditional diffusion model, we employ OpenCLIP ViT-H/14  [^2] https://github.com/mlfoundations/open_clip as the CLIP image encoder. This model’s transformer consists of 20 transformer blocks, each with a width of 2,048. The model is trained for 100k iterations with a batch size of 256, using a cosine noising schedule (Nichol & Dhariwal, 2021) with 1000 timesteps. For the inpainting and refining models, we use DINOv2-G/14  [^3] https://github.com/facebookresearch/dinov2 as the image encoder. We leverage the pretrained stable diffusion V2.1  [^4] https://huggingface.co/stabilityai/stable-diffusion-2-1-base, modifying the first convolution layer to accommodate additional conditions. These models are trained for 200k and 100k iterations, respectively, each with a batch size of 128, and a linear noise schedule with 1000 timesteps is applied. In the inference stage, we use the DDIM sampler with 20 steps and set the guidance scale w to 2.0 for PCDMs on all stages.
@@ Appendix C Additional Results
Figure 10: More examples show the refining conditional diffusion model on PIDM. It is worth noting that ours denotes results obtained only using the prior conditional diffusion model and inpainting conditional diffusion mode.
@@
C.1 More qualitative results with refining conditional diffusion model.
We show the qualitative results of PIDM combined with refining conditional diffusion model in Figure 10.
C.2 More Qualitative Comparisons for PCDMs
We provide additional examples for comparison with the state-of-the-art (SOTA) methods in Figure 11.
C.3 More qualitative ablation results on the DeepFashion dataset
@@
C.4 Random Samples
We show random samples from our proposed PCDMs in Figure 13 and Figure 14. The generated results are highly stable and realistic.
C.5 Impact of Pretrained
We show random samples from our proposed PCDMs in Figure 15. The results show that even without pretrained, the outcomes are still remarkably realistic. This indicates that the improvement in text reconstruction is primarily due to the rationality and superiority of our framework.
Figure 13: Random samples from DeepFashion dataset by PCDMs.
Figure 14: Random samples from Market-1501 dataset by PCDMs.
Figure 15: Compared with w/o pretrained weghts on PCDMs.
C.6 User Study
@@
C.7 Visualize the Diffusion Process
To facilitate a better understanding of the diffusion process, we have added visualizations for the second and third stages of the diffusion process, as shown in Figure 18 and Figure 19.
@@ C.7 Visualize the Diffusion Process
Figure 18: Visualize the diffusion process on the second stage.
Figure 19: Visualize the diffusion process on the third stage.